1. A bicycle is listed in a wholesaler's catalog at P5400. The bicycle is sold to a retailer for a net price of P4087.50. What rate trade discount was given?



2. A swimsuit with a list price of P3499 was sold to a retailer for P2624.25. What rate of trade discount is given?



3. A car accessory listed at P1136 has a net price of P760. What is the rate of discount?



4. Utik utik distributors offers discounts of 25%, 12.5%, and 4% on a line of products. For how much should an item be listed if it is to be sold for P113.40?



5. The net price of a product is P63.31. What is the suggested retail price (list price) of a discount off 35% was allowed?



6. In 1995, david bought 200m² lot for P200,000. This year, he is planning to sell the lot for P2,000,000. How much is the gross margin? What is the gross margin rate?




1
Expert's answer
2021-11-15T08:28:17-0500

Net Price = List Price -Trade Discount

Trade-Discount Rate = Trade Discount/List Price


1.

Trade Discount = "5400-4087.5=P1312.5"

Trade-Discount Rate = "1312.5/5400=0.24=24\%"


2.

Trade Discount = "3499-2624.25=P874.75"

Trade-Discount Rate = "874.75/3499=0.25=25\%"


3.

Trade Discount = "1136-760=P376"

Trade-Discount Rate = "376/1136=0.33=33\%"


4.

List Price = Net Price +Trade Discount = Net Price + Trade-Discount Rate x List Price

List Price = Net Price/(1 - Trade-Discount Rate)


for 25%:

List Price = "113.4\/(1-0.25)=P151.20"


for 12.5%:

List Price = "113.4\/(1-0.125)=P129.60"


for 4%:

List Price = "113.4\/(1-0.04)=P118.13"


5.

List Price = "63.31\/(1-0.35)=P97.40"


6.

Gross Margin = Revenue – Cost of Goods Bought

Gross Margin = "2000000-200000=P1800000"


Gross Margin Rate = (Revenue – Cost of Goods Bought) / Revenue

Gross Margin Rate = (2000000-200000)/2000000=0.9=90\%